OpenStudy (anonymous):

A farmer builds a pen with a perimeter of 116 feet. If the length is 5 feet more than the width , what is the width

OpenStudy (anonymous):

4w+10=116 4w=116-10 w=106/4 w=26.5 Now check 26.5*2 +(26.5+5)*2 =53+31.5*2 =53+63 =116